If f(x) = 2x3 – 3x2 + 7x + 1, what is the value of f[1, 2, 3, 4]?


1
Expert's answer
2021-07-13T05:10:46-0400
"f(1)=2\times1^3-3\times1^2+7\times1+1=7""f(2)=2\times2^3-3\times2^2+7\times2+1=19"

"f(3)=2\times3^3-3\times3^2+7\times3+1=49"

"f(4)=2\times4^3-3\times4^2+7\times4+1=109"